OpenStudy (anonymous):

Without drawing the graph of the equation, determine how many points the parabola has in common with the x-axis and whether its vertex lies above, on, or below the x-axis. y = –2x2 + x + 3 A. no points in common; vertex below x-axis B. 2 points in common; vertex above x-axis C. 2 points in common; vertex below x-axis D. 1 point in common; vertex on x-axis

OpenStudy (whpalmer4):

First: does this parabola open up or down?

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

@Bookworm14 @rose21 @undeadknight26

OpenStudy (whpalmer4):

You can tell just by looking at the coefficient of the \(x^2\) term. If it is positive, then for sufficiently large values of \(x\), \(y\) will be an increasingly large positive number. If it is negative, then again, as \(x\) gets farther away from the y-axis, \(y\) will again be an increasingly large positive number. In general, if you write the formula in the form \[y = ax^2+bx+c\]then if \(a>0\), the parabola opens upward, like a bowl. If \(a< 0\) the parabola looks like an inverted bowl, and opens downward.

OpenStudy (whpalmer4):

Next, you can use the discriminant of the quadratic equation to determine whether the parabola has 0, 1, or 2 points in common with the x-axis. The discriminant, \(\Delta\), is given by \(\Delta = b^2-4ac\), which you may recognize as the quantity under the square root symbol in the formula for the solutions to a quadratic equation. If \(\Delta=0\), you have a perfect square, and will have just one solution. If \(\Delta > 0\), you have two real solutions. If \(\Delta < 0\), you have two complex solutions, in a conjugate pair \((a\pm bi)\) A real solution occurs where \(y =0\). Complex solutions don't have an obvious appearance on the graph, unlike real solutions. A parabola which has only complex solutions appears suspended in midair above the x-axis (if opening upward) or hanging below the x-axis (if opening downward).

OpenStudy (whpalmer4):

One more bit of parabola lore: If you write the parabola in the standard quadratic form \(y=ax^2+bx+c\) then you can find the x-coordinate of the vertex with the simple equation \[x = -\frac{b}{2a}\]and then find the \(y\) value by plugging in that value of \(x\).